The perimeter of a regular hexagon is 3.04cm less than the perimeter of a regular pentagon. The perimeter of the regular hexagon is 21.06cm. What is the side length of the regular pentagon?
1
Expert's answer
2021-11-01T19:07:26-0400

Perimeter of the regular pentagon: "21.06+3.04=24.1" cm.

Side length of the regular pentagon: "l=\frac{P}{5}=\frac{24.1}{5}=4.82" cm.